How much water is needed to dilute 30. mL of 5.0 M sugar solution to 0.40 M?

410 mL of water is needed

no water is needed

350 mL of water is needed

380 mL of water is needed
Answer #1

Initial volume = 30mL

Molarity = 5.0M

So moles of sugar = molarity * Volume = 5.0M * 30 mL = 150 mmol

Lets say x mL of water is added

Total volume = (30+x)

So new molarity = Moles of sugar / New volume

Given,

new molarity = 0.40M

SO, 0.40M = 150mmol/(30+x)

(30+x ) = (150/0.40) mL

30+x = 375mL

x = 345 mL

So, the most approximate answer is 350 mL of water needs to be added.